OpenStudy (anonymous):

Paul plans to put concrete on a rectangular portion of his driveway. The portion is 12 feet long and 6 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Paul needs is $108.89. Which of the following is closest to the width of the portion of the driveway on which Paul plans to put concrete? [1 foot = 12inches; 1 yard = 3 feet] 3 feet 5 feet 8 feet 10 feet

OpenStudy (anonymous):

@johnweldon1993 (:

OpenStudy (johnweldon1993):

Well 12 feet= 4 yards 6 inches=0.166667 yards Rectangular portion area=4*0.166667*X yards (X is the unknown width) Paul needs a total cost of 108.89 and a cubic yard costs $98 So...how many cubic yards can he cover with that? =108.89/98= 1.1111 cubic yards Hence, 4*0.166667*X=1.1111 X=1.6666 yards which is equal to approximately 5 feet.

OpenStudy (anonymous):

thanks hun<3